### v1

@@ 5 MHD Simulations and the Radiation Pressure Dominated Thermal/Viscous Instability
As we discussed in the introduction, thermal and viscous instabilities were first theoretically predicted to exist in the inner, radiation pressure and electron scattering-dominated region of the Shakura & Sunyaev model [1] of black hole accretion discs [2, 3]. Standard accretion disc theory with radial advective cooling predicts an S-curve shape in the local accretion rate - surface density plane, with an upper stable slim disc branch and a lower stable gas pressure dominated branch [21]. One would therefore expect limit cycle behavior just as in the hydrogen or helium ionization instabilities in cataclysmic variables and X-ray binaries, and indeed such outbursting behavior has been predicted [177, 178]. However, observational evidence for the existence of radiation pressure dominated instability remains highly ambiguous. Complex classes of flaring behavior are definitely present in black holes and neutron stars, e.g. GRS 1915+105 [179], IGR J17091-3624 [180], the rapid burster MXB 1730-335 [181], V404 Cyg [182], the bursting pulsar GRO J1744-28 [183], the ULX 4XMM J111816.0-324910 in NGC 3621 [184], and Swift J1858.6-0814 [185]. Where masses and distances are known, these all appear to be accreting at or above the Eddington limit, and it has been proposed that at least some of this flaring behavior is associated with some version of the radiation pressure dominated instability [7, 186, 187, 188, 189, 190]. However, there are also many black hole X-ray binaries that achieve high Eddington ratios in the soft state, whose spectra are well-fit by standard geometrically thin discs, and that show negligible variability in their soft X-ray emission [26]. Why, then, are they so stable?
@@
A possible answer arose from early, vertically stratified shearing box simulations of MRI turbulence that incorporated diffusive radiation transport in the radiation pressure and electron scattering dominated regime [191]. Surprisingly, these simulations were able to achieve steady thermal equilibria, even though the time-averaged stress appeared to approximately scale with total thermal pressure rather than having a dependency on the much smaller gas pressure [192]. A cross-correlation analysis showed that the fluctuations in the thermal energy density responded to fluctuations in the turbulent energy with a lag of 5-15 local orbital periods, approximately the thermal time for these vertically stratified simulations. Dissipation of turbulence into thermal energy requires some time, in contrast to the instantaneous relation that is assumed in the standard alpha prescription (1), and this was proposed as an explanation for thermal stability [191]. The time-averaged stress still had an inverse relationship to surface density, however, so viscous instability was still a possibility [192].
@@
Unfortunately, this partial solution to the observed stability of the soft state was spurious. Later attempts to replicate the simulated thermal stability failed, starting with vertically stratified shearing boxes [193]. Even though the same stress to pressure time delay was present, these simulations were unable to maintain thermal equilibrium and always eventually underwent thermal runaway. The simulated thermal stability in [191] was due in part to the use of flux-limited diffusion as a radiation transport algorithm and to a relatively small horizontal box size. It is now known that such small horizontal domain sizes can dramatically affect the properties of MRI turbulence, and that the stresses have substantial large scale horizontal correlations [144]. Moreover, a scaling of stress with thermal pressure, which is what drives the thermal instability in the radiation pressure dominated regime, is only observed in gas pressure dominated shearing box simulations if the box size is larger than the thermal pressure scale height [105]. Runaway cooling in the radiation pressure dominated regime has also been observed in global general relativistic radiation MHD simulations [194, 195, 196].
@@
That turbulent fluctuations lead pressure fluctuations due to dissipation is not surprising on short time scales, but if the alpha prescription is to have any validity in a causal sense, then presumably pressure fluctuations on longer time scales alter the saturation level of turbulence [166, 105], and may even lead to stress fluctuations that lag pressure fluctuations [197]. Indeed, such a lag of order five orbital periods has been measured in vertically unstratified simulations with an optically thin cooling function that is artificially made to vary in time [198]. Strangely, this is smaller than the lag of pressure with respect to stress measured in optically thick vertically stratified simulations with radiation transport [191], but this may be because the optically thin cooling in [198] responds instantaneously to changes in pressure, unlike diffusive cooling. Measuring lags and leads as a function of fluctuation frequency, which can be done using Fourier methods [199], might be fruitful in radiation MHD simulations. The effects of simplified, fixed time delays on thermal stability have been investigated analytically by a number of authors. A lag of stress relative to pressure can reduce the thermal instability growth rate, but not stabilize it [200], whereas a lag of pressure relative to stress can be stabilizing [201]. The stochasticity of turbulent fluctuations themselves could in principle stabilize the disc if the fluctuations are large enough [202], but simulations of MRI turbulence in a nominally thermally unstable situation find that, while thermal instability can be slowed, it is still generally present [197].
@@
As we mentioned in §1, one possible explanation for thermal stability is that the inner parts of black hole accretion discs are supported against the vertical tidal gravity by magnetic, not thermal, pressure [13, 14]. Indeed, simulations of cooling from a hot optically thin flow (the hard state) to a cooler flow (the soft state) found a stable, long-lived magnetically supported flow [125]. Global radiation MHD simulations of discs supported by magnetic pressure are found not to undergo thermal runaways [194], and discs that are initialized with magnetic field topologies that are able to achieve a magnetically dominated state also avoid thermal runaways [203, 204, 196]. Stabilization can also be achieved even if the midplane regions are supported by thermal pressure, but most of the accretion occurs at altitude in a magnetically dominated region: radiation MHD simulations of the magnetically elevated disc flows discussed above in §3.2 find that a radiation pressure dominated midplane can be stabilized in this regime [205]. Such discs fit neither of the two categories of thin or slim, and the authors refer to them as “puffy" discs. Spectra of such discs have been computed in the mildly sub-Eddington regime, and resemble thermal spectra with a warm corona [206].
@@
But this begs the question as to whether magnetically supported discs can have spectra that truly resemble the non-variable, strongly thermal spectra with no energetically significant corona that can exist in the high/soft state of black hole X-ray binaries [25]. Can a magnetically supported disc produce such thermal spectra, and can they show insignificant variability [26]? Note that observations by the Imagning X-ray Polarimetry Explorer can rule out the presence of large scale coherent magnetic fields due to Faraday rotation [207], which may be another way of constraining this solution to the observed lack of thermal instability in the high/soft state.
@@
Yet another possible solution to the lack of observational evidence of thermal instability in the soft state is that modern simulations have not yet accurately captured the saturation behavior of the MRI under radiation pressure dominated conditions. Indeed, it has only recently been convincingly demonstrated that MRI turbulence under incompressible conditions has an inertial range with constant energy flux from large scales to small scales [96]. The simulation employed was a vertically unstratified shearing box, and even this required considerable computational resources. The vertically stratified and global radiation MHD simulations that find thermal instability have nowhere near the dynamic range to explore this issue. On the other hand, the photon diffusion scale is likely not much smaller than the driving scale in a radiation pressure dominated accretion disc [106], so that compressive damping is likely important on large scales, but incompressible fluctuations may still drive a cascade to smaller scales. It may perhaps be that the saturation properties of MRI turbulence in radiation pressure dominated environments still requires a high dynamic range, and it might be worthwhile expending computational resources to investigate this problem.
@@
While our discussion here has focused on X-ray binaries, accretion discs in luminous active galactic nuclei (AGN) are expected to be even more radiation pressure dominated, so one might ask how thermal/viscous instabilities might manifest there. A major complication is that the fiducial temperatures are much lower (ultraviolet) in AGN compared to X-ray binaries, and simply on the basis of the free-free opacity therefore being larger, AGN discs were predicted early on to be thermally stable [9]. However, vertically stratified shearing box simulations with just free-free and electron scattering opacities are apparently unable to achieve long-lived thermal equilibria [208]. On the other hand, additional opacity sources are also present. AGN discs have densities and temperatures similar to those in the envelopes of massive stars [208], and iron can significantly enhance the Rosseland mean opacity $\kappa$ over that of pure electron scattering $\kappa_{\mathrm{T}}$ near temperatures $\sim 2\times 10^{5}$ K [209]. Vertically stratified shearing box simulations are able to achieve long-lived thermal equilibria when this opacity enhancement is present [208]. In the standard model, where radiation pressure provides hydrostatic support against the tidal gravity, the iron opacity region is located at radius given by
@@
Global simulations in this radial range find, indeed, that thermal runaways do not occur in this radial range in radiation pressure dominated flows [116]. However, considerable variability is still present. The iron opacity peak causes vertical density inversions which are convectively unstable, and this leads to convective/radiative cyclic behavior with variable enhancements of MRI turbulent stresses. These in turn cause radial concentrations of surface mass density that appear and disappear. Strong luminosity variability is associated with these cycles, with a characteristic time scale of order a year, consistent with the thermal time. This might therefore be a possible explanation for the characteristic time scale inferred in damped random walk models of AGN optical variability [210].
Away from the iron opacity dominated region, thermal stability can again be achieved if the disc is supported against tidal gravity by magnetic pressure [211, 212]. In addition to coherent and turbulent Maxwell stresses, one such simulation had substantial radiation viscosity: the off-diagonal terms in the radiation stress tensor in the optically thin layers drove more than half the accretion rate [211].
As noted in §3.2, a recent cosmological zoom-in simulation has produced a very strongly magnetized accretion disc on scales of order 100 gravitational radii [128, 127]. Such a disc is much less dense and less gravitationally unstable than a standard Shakura-Sunyaev disc. Analytic scalings based on the simulation results [129] suggest that radiation pressure will become more important near the black hole at near or super-Eddington accretion rates. A very recent extension of these simulations to several gravitational radii around the black hole does indeed find that radiation pressure is comparable, but does not dominate, magnetic and turbulent pressure [213]. These simulations generally have much higher inflow speeds and lower surface mass densities and vertical optical depths than both standard Shakura-Sunyaev models [129] and magnetically elevated discs that are initialized with tori around the black hole. Whether this is consistent with observational constraints remains to be seen, but it appears likely that there remains a large parameter space of strongly magnetized accretion discs that has yet to be explored.
Finally, before leaving this section we should point out that there are also now many global simulations of super-Eddington accretion onto black holes and neutron stars [111, 214, 215, 216, 217], which may be of relevance to the highest luminosity sources, e.g. ultra-luminous X-ray sources. Thermal stability is less of an issue in this regime, and instead the focus has been on the efficacy of photon trapping and radial advection of heat, compared to vertical diffusive and advective escape. Outflows driven by continuum radiation pressure are also an important issue here. We refer the interested reader to the cited references for more information.
@@ 6 Conclusion: Outstanding Questions and Directions for Future Research
    Thermal and viscous instabilities are most evident in outbursting cataclysmic variables and X-ray binaries, and these systems still provide the strongest constraints on the angular momentum transport mechanism that operates in accretion discs in this regime. As we discussed here, the $\alpha$-based disc instability model, with some ad hoc tweaks, can reproduce observed outbust light curves. Some aspects of these light curves can also be reproduced with models based on local simulations of MRI turbulence, but more work is needed to investigate and constrain magnetic angular momentum transport in these systems. In particular, simulations that allow for the possibility of magnetocentrifugal winds, and that also address the physics of the quiescent state, are needed. The simulations that do exist demonstrate that composition matters: hydrogen and helium discs are predicted to behave very differently, and more work is needed to see how this might manifest observationally.
@@
    More global simulations that incorporate the coupling between magnetic angular momentum transport mechanisms and the thermodynamics (heating and cooling) of the disc are needed, in all astrophysical contexts from cataclysmic variables to active galactic nuclei. There is a rich interplay that is only just beginning to be elucidated. Particularly important is how thermodynamics couples not only to weak field MRI turbulence but also to transport by large scale magnetic field stresses, including magnetocentrifugal winds.
@@
    At the same time, local shearing box experiments that are carefully constructed to address basic physics questions still need to be pursued. On what fluctuation time scales do stresses lead and lag thermal pressure fluctuations? How does field topology and magnetic Prandtl number affect turbulent stresses, especially when coupled with thermodynamics? Can the recent successful capture of an inertial range in incompressible MRI turbulence [96] be generalized to compressive MRI turbulence in a radiation pressure dominated plasma? How does thermodynamics couple to the dynamo that maintains a strongly magnetized turbulence state? While such local experiments neglect the radial coupling due to radial flows and large scale magnetic fields, answers to these fundamental questions would still be of great use.
@@
    Finally, in addition to theoretical work, more observational constraints on the presence and strength of turbulence in different astrophysical discs are desperately needed. In addition to, e.g., modeling light curves, useful information can be obtained by measuring the turbulent broadening of spectral lines in protoplanetary discs and AGN discs [218, 219, 220], or even in binaries with compact stars [221, 222]. In addition, imaging observations of rings in protoplanetary discs [223] provide constraints in that context. Direct constraints on the strength of magnetic fields in discs can also be obtained through detections of polarized disc emission [207].

### v2

@@ 5 MHD Simulations and the Radiation Pressure Dominated Thermal/Viscous Instability
As we discussed in the introduction, thermal and viscous instabilities were first theoretically predicted to exist in the inner, radiation pressure and electron scattering-dominated region of the Shakura & Sunyaev model (sha73) of black hole accretion discs [lig74, sha76]. Standard accretion disc theory with radial advective cooling predicts an S-curve shape in the local accretion rate - surface density plane, with an upper stable slim disc branch and a lower stable gas pressure dominated branch [che95]. One would therefore expect limit cycle behavior just as in the hydrogen or helium ionization instabilities in cataclysmic variables and X-ray binaries, and indeed such outbursting behavior has been predicted [hon91, szu98]. However, observational evidence for the existence of radiation pressure dominated instability remains highly ambiguous. Complex classes of flaring behavior are definitely present in black holes and neutron stars, e.g. GRS 1915+105 [bel00], IGR J17091-3624 [alt11], the rapid burster MXB 1730-335 [bag15], V404 Cyg [mot17], the bursting pulsar GRO J1744-28 [cou18], the ULX 4XMM J111816.0-324910 in NGC 3621 [mot20], and Swift J1858.6-0814 [vin23]. Where masses and distances are known, these all appear to be accreting at or above the Eddington limit, and it has been proposed that at least some of this flaring behavior is associated with some version of the radiation pressure dominated instability [taa84, can96, bel97, taa97, nay00, jan00]. However, there are also many black hole X-ray binaries that achieve high Eddington ratios in the soft state, whose spectra are well-fit by standard geometrically thin discs, and that show negligible variability in their soft X-ray emission [gie04]. Why, then, are they so stable?
@@
A possible answer arose from early, vertically stratified shearing box simulations of MRI turbulence that incorporated diffusive radiation transport in the radiation pressure and electron scattering dominated regime [hir09a]. Surprisingly, these simulations were able to achieve steady thermal equilibria, even though the time-averaged stress appeared to approximately scale with total thermal pressure rather than having a dependency on the much smaller gas pressure [hir09b]. A cross-correlation analysis showed that the fluctuations in the thermal energy density responded to fluctuations in the turbulent energy with a lag of 5-15 local orbital periods, approximately the thermal time for these vertically stratified simulations. Dissipation of turbulence into thermal energy requires some time, in contrast to the instantaneous relation that is assumed in the standard alpha prescription (1), and this was proposed as an explanation for thermal stability [hir09a]. The time-averaged stress still had an inverse relationship to surface density, however, so viscous instability was still a possibility [hir09b].
@@
Unfortunately, this partial solution to the observed stability of the soft state was spurious. Later attempts to replicate the simulated thermal stability failed, starting with vertically stratified shearing boxes [jia13b]. Even though the same stress to pressure time delay was present, these simulations were unable to maintain thermal equilibrium and always eventually underwent thermal runaway. The simulated thermal stability in hir09a was due in part to the use of flux-limited diffusion as a radiation transport algorithm and to a relatively small horizontal box size. It is now known that such small horizontal domain sizes can dramatically affect the properties of MRI turbulence, and that the stresses have substantial large scale horizontal correlations [sim12]. Moreover, a scaling of stress with thermal pressure, which is what drives the thermal instability in the radiation pressure dominated regime, is only observed in gas pressure dominated shearing box simulations if the box size is larger than the thermal pressure scale height [ros16]. Runaway cooling in the radiation pressure dominated regime has also been observed in global general relativistic radiation MHD simulations [sad16, mis16, lis22].
@@
That turbulent fluctuations lead pressure fluctuations due to dissipation is not surprising on short time scales, but if the alpha prescription is to have any validity in a causal sense, then presumably pressure fluctuations on longer time scales alter the saturation level of turbulence [lat12, ros16], and may even lead to stress fluctuations that lag pressure fluctuations [ros17]. Indeed, such a lag of order five orbital periods has been measured in vertically unstratified simulations with an optically thin cooling function that is artificially made to vary in time [hel22a]. Strangely, this is smaller than the lag of pressure with respect to stress measured in optically thick vertically stratified simulations with radiation transport [hir09a], but this may be because the optically thin cooling in hel22a responds instantaneously to changes in pressure, unlike diffusive cooling. Measuring lags and leads as a function of fluctuation frequency, which can be done using Fourier methods [now99], might be fruitful in radiation MHD simulations. The effects of simplified, fixed time delays on thermal stability have been investigated analytically by a number of authors. A lag of stress relative to pressure can reduce the thermal instability growth rate, but not stabilize it [lin11], whereas a lag of pressure relative to stress can be stabilizing [cie12]. The stochasticity of turbulent fluctuations themselves could in principle stabilize the disc if the fluctuations are large enough [jan12], but simulations of MRI turbulence in a nominally thermally unstable situation find that, while thermal instability can be slowed, it is still generally present [ros17].
@@
As we mentioned in §1, one possible explanation for thermal stability is that the inner parts of black hole accretion discs are supported against the vertical tidal gravity by magnetic, not thermal, pressure [beg07, oda09]. Indeed, simulations of cooling from a hot optically thin flow (the hard state) to a cooler flow (the soft state) found a stable, long-lived magnetically supported flow [mac06]. Global radiation MHD simulations of discs supported by magnetic pressure are found not to undergo thermal runaways [sad16], and discs that are initialized with magnetic field topologies that are able to achieve a magnetically dominated state also avoid thermal runaways [mor18, mis22, lis22]. Stabilization can also be achieved even if the midplane regions are supported by thermal pressure, but most of the accretion occurs at altitude in a magnetically dominated region: radiation MHD simulations of the magnetically elevated disc flows discussed above in §3.2 find that a radiation pressure dominated midplane can be stabilized in this regime [lan19]. Such discs fit neither of the two categories of thin or slim, and the authors refer to them as “puffy" discs. Spectra of such discs have been computed in the mildly sub-Eddington regime, and resemble thermal spectra with a warm corona [wie22].
@@
But this begs the question as to whether magnetically supported discs can have spectra that truly resemble the non-variable, strongly thermal spectra with no energetically significant corona that can exist in the high/soft state of black hole X-ray binaries [dav06]. Can a magnetically supported disc produce such thermal spectra, and can they show insignificant variability [gie04]? Note that observations by the Imaging X-ray Polarimetry Explorer might be able to constrain the presence of large scale coherent magnetic fields due to Faraday rotation [bar24], possibly providing a way of testing this solution to the observed lack of thermal instability in the high/soft state.
@@
Yet another possible solution to the lack of observational evidence of thermal instability in the soft state is that modern simulations have not yet accurately captured the saturation behavior of the MRI under radiation pressure dominated conditions. Indeed, it has only recently been convincingly demonstrated that MRI turbulence under incompressible conditions has an inertial range with constant energy flux from large scales to small scales [kaw24]. The simulation employed was a vertically unstratified shearing box, and even this required considerable computational resources. The vertically stratified and global radiation MHD simulations that find thermal instability have nowhere near the dynamic range to explore this issue. On the other hand, the photon diffusion scale is likely not much smaller than the driving scale in a radiation pressure dominated accretion disc [tur03], so that compressive damping is likely important on large scales, but incompressible fluctuations may still drive a cascade to smaller scales. It may perhaps be that the saturation properties of MRI turbulence in radiation pressure dominated environments still requires a high dynamic range, and it might be worthwhile expending computational resources to investigate this problem.
@@
While our discussion here has focused on X-ray binaries, accretion discs in luminous active galactic nuclei (AGN) are expected to be even more radiation pressure dominated, so one might ask how thermal/viscous instabilities might manifest there. A major complication is that the fiducial temperatures are much lower (ultraviolet) in AGN compared to X-ray binaries, and simply on the basis of the free-free opacity therefore being larger, AGN discs were predicted early on to be thermally stable [pri76]. However, vertically stratified shearing box simulations with just free-free and electron scattering opacities are apparently unable to achieve long-lived thermal equilibria [jia16]. On the other hand, additional opacity sources are also present. AGN discs have densities and temperatures similar to those in the envelopes of massive stars [jia16], and iron can significantly enhance the Rosseland mean opacity $\kappa$ over that of pure electron scattering $\kappa_{\mathrm{T}}$ near temperatures $\sim 2\times 10^{5}$ K [jia15]. Vertically stratified shearing box simulations are able to achieve long-lived thermal equilibria when this opacity enhancement is present [jia16]. In the standard model, where radiation pressure provides hydrostatic support against the vertical tidal gravity, the iron opacity region is located at radius given by
@@
Global simulations find, indeed, that thermal runaways do not occur in this radial range in radiation pressure dominated flows [jia20], see Fig. 4. However, considerable variability is still present. The iron opacity peak causes vertical density inversions which are convectively unstable, and this leads to convective/radiative cyclic behavior with variable enhancements of MRI turbulent stresses. These in turn cause radial concentrations of surface mass density that appear and disappear. Strong luminosity variability is associated with these cycles, with a characteristic time scale of order a year, consistent with the thermal time. This might therefore be a possible explanation for the characteristic time scale inferred in damped random walk models of AGN optical variability [bur21].
Away from the iron opacity dominated region, thermal stability can again be achieved if the disc is supported vertically by magnetic pressure [jia19, cur23]. In addition to coherent and turbulent Maxwell stresses, one such simulation had substantial radiation viscosity: the off-diagonal terms in the radiation stress tensor in the optically thin layers drove more than half the accretion rate [jia19].
As noted in §3.2, a recent cosmological zoom-in simulation has produced a very strongly magnetized accretion disc on scales of order 100 gravitational radii [hop24a, hop24b]. Such a disc is much less dense and less gravitationally unstable than a standard Shakura-Sunyaev disc. Analytic scalings based on the simulation results [hop24c] suggest that radiation pressure will become more important near the black hole at near or super-Eddington accretion rates. A very recent extension of these simulations to several gravitational radii around the black hole does indeed find that radiation pressure is comparable, but does not dominate, magnetic and turbulent pressure [hop25]. These simulations generally have much higher inflow speeds and lower surface mass densities and vertical optical depths than both standard Shakura-Sunyaev models [hop24c] and magnetically elevated discs that are initialized with tori around the black hole. Whether this is consistent with observational constraints remains to be seen, but it appears likely that there remains a large parameter space of strongly magnetized accretion discs that has yet to be explored.
Finally, before leaving this section we should point out that there are also now many global simulations of super-Eddington accretion onto black holes and neutron stars [jia14, sad14, mck14, sadnar16, ino24], which may be of relevance to the highest luminosity sources, e.g. ultra-luminous X-ray sources. Thermal stability is less of an issue in this regime, and instead the focus has been on the efficacy of photon trapping and radial advection of heat, compared to vertical diffusive and advective escape. Outflows driven by continuum radiation pressure are also an important issue here. We refer the interested reader to the cited references for more information.
@@ 6 Conclusion: Outstanding Questions and Directions for Future Research
    Thermal and viscous instabilities are most evident in outbursting cataclysmic variables and X-ray binaries, and these systems still provide the strongest constraints on the angular momentum transport mechanism that operates in accretion discs in this regime. As we discussed here, the $\alpha$-based disc instability model, with some ad hoc tweaks, can reproduce observed outburst light curves. Some aspects of these light curves can also be reproduced with models based on local simulations of MRI turbulence, but more work is needed to investigate and constrain magnetic angular momentum transport in these systems. In particular, simulations that allow for the possibility of magnetocentrifugal winds, and that also address the physics of the quiescent state, are needed. The simulations that do exist demonstrate that composition matters: hydrogen and helium discs are predicted to behave very differently, and more work is needed to see how this might manifest observationally.
@@
    More global simulations that incorporate the coupling between magnetic angular momentum transport mechanisms and the thermodynamics (heating and cooling) of the disc are needed, in all astrophysical contexts from cataclysmic variables to active galactic nuclei. There is a rich interplay that is only just beginning to be elucidated. Particularly important is how thermodynamics couples not only to weak field MRI turbulence but also to transport by large scale magnetic field stresses, including magnetocentrifugal winds. How angular momentum transport is coupled to thermodynamics is also extremely important in understanding the propagation of heating and cooling fronts in the disk during the dwarf nova outburst cycle, though this is likely to be very challenging given current computational resources.
@@
    At the same time, local shearing box experiments that are carefully constructed to address basic physics questions still need to be pursued. On what fluctuation time scales do stresses lead and lag thermal pressure fluctuations? How does field topology and magnetic Prandtl number affect turbulent stresses, especially when coupled with thermodynamics? Can the recent successful capture of an inertial range in incompressible MRI turbulence [kaw24] be generalized to compressive MRI turbulence in a radiation pressure dominated plasma? How does thermodynamics couple to the dynamo that maintains a strongly magnetized turbulence state? While such local experiments neglect the radial coupling due to radial flows and large scale magnetic fields, answers to these fundamental questions would still be of great use.
@@
    Finally, in addition to theoretical work, more observational constraints on the presence and strength of turbulence in different astrophysical discs are desperately needed. In addition to, e.g., modeling light curves, useful information can be obtained by measuring the turbulent broadening of spectral lines in protoplanetary discs and AGN discs [Flaherty+2018, Flaherty+2024, Ochman+2024], or even in binaries with compact stars [Horne1995, Skidmore+2000]. In addition, imaging observations of rings in protoplanetary discs [Rosotti+2020] provide constraints in that context. Direct constraints on the strength of magnetic fields in discs can also be obtained through detections of polarized disc emission [bar24].